Find the amount at the end of 5 years if R17 535 is invested at 2.45% compounded quarterly.

1
Expert's answer
2021-11-25T20:36:01-0500

compounded quarterly,n=4


"A=P(1+\frac{r}{n})^{nt}"


"=R17,535(1+\frac{0.0245}{4})^{4×5}"


"=17,535(1.006125)^{20}"


"=R19812.74"